The energy difference between two electronic states is 399.1kj /mol .Calculate the wavelenth and frequency of light emitted when an electron drops form a higher to a lower state .(Plank`s constant =6.6262 ×10^-34 Js).

Grade:12

1 Answers

Vikas TU
7 years ago
Use the Planck’s energy equation,
E = hv 
399.1 x 10^3 = 6.6262 x 10^-34 x v
v = 6579 x 10^30
and for the wavelength calculation,
v = c/k
k = c/v => (3*10^8)/(6579*10^30) => 4.55 x 10^-26 m